# Train-before-Test Harmonizes Language Model Rankings

### v1

@@ A.3 Evaluation Setup
For our train-before-test evaluations, we fine-tune each model for five epochs and select the best-performing checkpoint based on evaluations on a separate validation set. To reduce memory consumption, we employ parameter-efficient fine-tuning (PEFT) [34, 47], allowing us to conduct most experiments on Quadro RTX 6000, Tesla V100-SXM2-32GB and NVIDIA A100-SXM4-80GB GPUs.
We use fixed hyperparameters across all experiments. These hyperparameters were not tuned for any particular model or benchmark. Specifically, we use AdamW with a learning rate of 2e-5 and weight decay of 0.01. For PEFT, we use a LoRA configuration with rank 8, $\alpha=32$, and dropout 0.1.
For models that do not exhibit performance improvements after fine-tuning, we simply report their pre-fine-tuning results. Since all training datasets considered in our work are publicly available, models may have already trained on them during pre-training. In certain cases, this may lead to further fine-tuning not being beneficial.
@@
For instruction-tuned models, we assess performance both with and without applying chat templates and use the best performing one under direct evaluation. In the train-before-test setting, we similarly fine-tune two variants: one where the training data are formatted with chat templates and one without. We similarly select the best performing approach. We observe that not using a chat template often leads to better performance than using a chat template, both for direct evaluation and train-before-test.
@@ B.2 Perplexity Ranking Agreement
We exclude Gemma models from our perplexity agreement experiments, as lm-eval-harness provides unreliable perplexity measurements for Gemma models333 See discussion at https://github.com/huggingface/transformers/issues/29250. . We report the bits per byte (BPB) for the Gemma models in Table 3. While the BPB values for Gemma on arXiv (the dataset with the shortest average sequence length) are mostly reasonable, the performance on StackExchange and Wikipedia is notably worse, even compared to smaller models like Pythia-410M.
@@ B.3 PC1 Score under Train-before-Test
[…]
@@ C.2 Downstream Ranking Agreement
We reproduce the ranking correlation figures of the main text using the modified Kendall’s $\tau$ which treats non-statistically significant performance differences as ties. See Figure 12 and 13; as well as Figure 14 and Figure 15 for more detailed results. We observe that accounting for statistical significance in models’ performance differences leads to slightly higher ranking correlations, as measured by Kendall’s $\tau$-b. For direct evaluation, average agreement increases from 0.51 to 0.57. For train-before-test, average agreement increases from 0.74 to 0.77. Therefore, train-before-test continues to lead to large improvements in raking agreement (from Kendall’s $\tau$-b 0.57 to 0.77).
@@
rank1[$m$] $=$ position of $m$ in $order_{1}$ (starting from 1)
rank2[$m$] $=$ position of $m$ in $order_{2}$ (starting from 1)
@@
Figure 12: Mean ranking agreement between each benchmark and all others, measured by Kendall’s $tau$-b, with non-statistically significant performance differences being treated as ties. We calculate Kendall’s $\tau$-b between each benchmark and every other one, and then average. Compared to direct evaluation, train-before-test consistently improves ranking agreement–often by a large margin. On average, the overall average Kendall’s $\tau$ is 0.57 for direct evaluation and 0.77 for train-before-test.
@@
Figure 13: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right), measured by Kendall’s $tau$-b, with non-statistically significant performance differences being treated as ties. We consider language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med) categories. Kendall’s $\tau$-b is averaged across all pairs of benchmarks that belong to two given categories. The diagonal represents the intra-category agreement and the others represent the inter-category agreement. train-before-test improves both intra- and inter-category ranking agreement in all instances.
Figure 14: Cross benchmark ranking agreement under direct evaluation, measured by Kendall’s $tau$-b with insignificant model comparisons treated as ties.
Figure 15: Cross benchmark ranking agreement under train-before-test, measured by Kendall’s $tau$-b with insignificant model comparisons treated as ties.
Appendix D Broader Impacts and Limitations
Due to resource constraints, we do not conduct exhaustive hyperparameter searches for each model and benchmark. Instead, we use the same set of hyperparameters for all models and benchmarks. We do not tune these hyperparameters for any given model or benchmark. Therefore, all models received the same amount of hyperparameter search budget, albeit small.
We restrict our analysis to benchmarks that have a train set. However, recent benchmarks seldom include a train split. Thus, most of the benchmarks considered in our analysis, while typically highly cited and influential, were released before 2022.
We do not anticipate any direct societal impacts from this work, such as potential malicious or unintended uses, nor do we foresee any significant concerns involving fairness, privacy, or security considerations. Additionally, we have not identified potential harms resulting from the application of this technology.

### v2

@@ A.2 Evaluation Setup
For our train-before-test evaluations, we fine-tune each model for five epochs and select the best-performing checkpoint based on evaluations on a separate validation set. We use the AdamW optimizer with a weight decay of 0.01. For each model-benchmark combination, we perform a hyperparameter search over three learning rates $\{1e-5,2e-5,5e-5\}$ and select the optimal one based on validation performance. To reduce memory consumption, we employ parameter-efficient fine-tuning (PEFT) [38, 55], We use a LoRA configuration with rank 8, $\alpha=32$, and dropout 0.1. Most of our experiments are conducted on Quadro RTX 6000, Tesla V100-SXM2-32GB and NVIDIA A100-SXM4-80GB GPUs.
In cases where models show no performance improvement after fine-tuning, we report their pre-fine-tuning results. This scenario is rare and typically occur with smaller models (less than 500M parameters) that lack the capacity to perform certain tasks, resulting in near-random performance both before and after fine-tuning. Additionally, since all training datasets in our study are publicly available, some models may have encountered this data during pre-training, potentially limiting the benefits of additional fine-tuning.
@@
For instruction-tuned models, we evaluate performance both with and without chat templates, selecting the configuration that yields better results. Specifically, during direct evaluation, we assess model performance on the validation set under both conditions and apply the better-performing configuration to the test set. In the train-before-test setting, we similarly fine-tune two variants: one with training data formatted using chat templates and one without. We then select the approach that achieves the best performance on the validation set for final evaluation.
@@ B.2 Perplexity Ranking Agreement
We exclude Gemma models from our perplexity agreement experiments, as lm-eval-harness provides unreliable perplexity measurements for Gemma models [^3] See discussion at https://github.com/huggingface/transformers/issues/29250. . We report the bits per byte (BPB) for the Gemma models in Table 3. While the BPB values for Gemma on arXiv (the dataset with the shortest average sequence length) are mostly reasonable, the performance on StackExchange and Wikipedia is notably worse, even compared to smaller models like Pythia-410M.
@@ B.3 PC1 Score under Train-before-Test
Table 4: The models used in Figure 7. The number of training tokens of these models is publicly available. We compute the number of pre-training FLOPs as $6\times\text{\#Parameters}\times\text{\#Tokens}$.
@@ C.2 Downstream Ranking Agreement
We reproduce the ranking correlation figures of the main text using the modified Kendall’s $\tau$ which treats non-statistically significant performance differences as ties. See Figure 12 and 13; as well as Figure 14 and Figure 15 for more detailed results. We observe that accounting for statistical significance in models’ performance differences leads to slightly higher ranking correlations, as measured by Kendall’s $\tau$-b. For direct evaluation, average agreement increases from 0.52 to 0.58. For train-before-test, average agreement increases from 0.76 to 0.77. Therefore, train-before-test continues to lead to large improvements in raking agreement (from Kendall’s $\tau$-b 0.58 to 0.77).
@@
rank1[$m$] == position of $m$ in $order_{1}$ (starting from 1)
rank2[$m$] == position of $m$ in $order_{2}$ (starting from 1)
@@
Figure 12: Mean ranking agreement between each benchmark and all others, measured by Kendall’s tautau-b, with non-statistically significant performance differences being treated as ties. We calculate Kendall’s $\tau$-b between each benchmark and every other one, and then average. Compared to direct evaluation, train-before-test consistently improves ranking agreement–often by a large margin. On average, the overall average Kendall’s $\tau$ is 0.58 for direct evaluation and 0.77 for train-before-test.
@@
Figure 13: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right), measured by Kendall’s tautau-b, with non-statistically significant performance differences being treated as ties. We consider language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med) categories. Kendall’s $\tau$-b is averaged across all pairs of benchmarks that belong to two given categories. The diagonal represents the intra-category agreement and the others represent the inter-category agreement. train-before-test improves both intra- and inter-category ranking agreement in all instances.
Figure 14: Cross benchmark ranking agreement under direct evaluation, measured by Kendall’s tautau-b with insignificant model comparisons treated as ties.
Figure 15: Cross benchmark ranking agreement under train-before-test, measured by Kendall’s tautau-b with insignificant model comparisons treated as ties.